OpenStudy (anonymous):

Jason and Kyra have rectangular-shaped vegetable gardens. The width of Kyra’s garden is three times the width of Jason’s garden. The area of Kyra’s garden is equal to the area of Jason’s garden. Which statement is true? Kyra’s garden has the same perimeter as Jason’s garden. The length of Kyra’s garden is one-third of the length of Jason’s garden. The length of Kyra’s garden is three times the length of Jason’s garden. Kyra’s garden has the same length as Jason’s garden. one year ago

OpenStudy (anonymous):

@AkashdeepDeb @jim_thompson5910

OpenStudy (akashdeepdeb):

Let both of their areas be A Let Jason's gardens' width be x And so Kyra's garden width = 3x So Jason length = A/x ' Kyra length = A/3x Find ratio of Jason length: Kyra length = A/x / A/3x 3:1 Which one do you think'd be correct and why?

OpenStudy (anonymous):

Third or second, but I think third

OpenStudy (akashdeepdeb):

It is actually the second one good try though! See because, if the have the same AREA but kyra had MORE breadth then obviously Jason would have MORE length to compensate for the EQUAL AREA! Getting me? :)

OpenStudy (anonymous):

Yeah thank!!!!

OpenStudy (akashdeepdeb):

:)